a winery has a barrel to hold chardonnay. An inlet pipe can fill the barrel in 11 hours, while an outlet pipe can empty the barrel in 14 hours. how long will it take to fill the barrel if both the inlet and outlet poprs are open
Answer by nerdybill(7384)

a winery has a barrel to hold chardonnay. An inlet pipe can fill the barrel in 11 hours, while an outlet pipe can empty the barrel in 14 hours. how long will it take to fill the barrel if both the inlet and outlet poprs are open
.
Let x = time (hours) to fill barrel
then
x(1/11 - 1/14) = 1
multiplying both sides by (11)(14):
x(14 - 11) = (11)(14)
x(3) = 154
x = 154/3
x = 51.33 hours